### v4

# Measurement-induced physics and stochastic dynamics in systems of single-body quasiparticle and many-body microscopic cases

Abstract
In this article, we focus on measurement-induced physics and stochastic dynamics in the systems of single-body quasiparticle and many-body microscopic cases. In the former case, it shows a smooth evolution from a non-Gaussian (Poission or binomial distribution) with local dephasing to a Gaussian with global dephasing (quantum microscopic to classical macroscopic) where in late time there is a global symmetry that consists of exponentially many pure states due to the suppressed (stochastic) fluctuation and heating; In the latter case, it shows exponentially many distinct (Krylov) sectors due to the fragmentation and dynamical constrain on Hilbert-space. A gapless-to-gapped transition ((sub)volume law to area-law) during evolution is evident for the former case, where, however, the asymptotic approaching area-law scaling (system-size-independence) even appears in early time with the increasing system-size toward infinity. A novel purifying lasting for a short time is found in the early stage of the former case, during which there is long-range power law decay and scrambling before the Lieb-Robinson bound. Meanwhile, in the latter case, local non-Gaussian interactions (with a large power-law exponent) are present, maintaining an invariant gapless feature. Except a comprehensive study for the both cases, our scheme avoid the interaction effect as well as the selections on measured states due to the measurement apparatus, and thus easier evaluatin on the expectations of non-Hermitian observables.
I Introduction
It has been studied that a pair of coupled identical Sachdev-Ye-Kitaev (SYK) models with complex fermions provide globally chaotic systems which allowing the quantum mutual information to traverse through the wormhole[1], where the robust ergodicity against the localization is supported by the revival oscillations between the two chaotic subsystems[1, 2]. Similarly, a periodic initial state revivals in a many-body system indicate the nonthermal feature in the quench dynamics up to long-time limit (infinite temperature or high energy density)[142, 172, 22], which is distinct to the conventional many-body scars. The non-Hermitian dynamics can arises from the quantum dissipation or continuous measurement together with postselected trajectories (exponentially costly[158, 28]). The non-Hermitian skin effect with finite dissipation (imaginary) gap in complex spectrum shows area-law entanglement scaling. The skin effect is caused by the nonreciprocal dissipation and thus suppressing the entanglement diffusion. Despite sharing the boundary localization, the edge burst[57] additionally require closing dissipative gap[28, 53, 54], origin from a universal bulk-edge scaling relation. Moreover, the multi-partite entanglement can be effectively manipulated through the non-Hermitian (asymmetry) coupling. Note that in this article, for skin effect, we only refer to the nonreciprocal one, i.e., excluding the reciprocal induced by $Z_{2}$-type symmetry[61, 64], and for coherence, we only refer to the one happen in non-Markov case, such that the coherence field (pattern) corresponding to a closed system in a pure state, instead of the global conservation (topological invariance). In non-Hermitian (open) system, there is unidirectional disentangling due to the coupling to the external disorder. Then there is possibly be infinite number of bound states (in gapped regime) discretically, among which are mutually uncorrelated and the level statistics follows Gaussian ensemble (random matrices). The only way to reveal such external disorder (or non-Hermitian perturbation) is to continuously monitor a individual quantum trajectory, which is a separable pure state (corresponding to a rank-1 subspace) with the continuously measured outcomes be postselected. Here the "individuality" is guaranteed by conditioning the evolution to the absence of quantum jump (sudden change of state) as well as the exceptional point[12, 13, 14, 15]. Such individual trajectory subject to unitary evolution (by certain Hermitian operator $H$) $|\psi(t)\rangle=\frac{e^{-iHt}|\psi_{0}\rangle}{||e^{-iHt}|\psi_{0}\rangle||}$ is conditioned during the continuously measurement and thus support the long-range entanglement, while the repeated local measurement (monitoring) will suppress the short-range entanglement[44, 39] by alternating the random unitaries and introducing the quantum jump processes accompanied by the density fluctuation.
In this article, instead of exploring the measurement-induced transition, we study the two extreme cases, the limit of weak but continuous measurement and strong but sparse measurement, such that the measurement apparatus plays little role in our scheme. This also provides great convenience in measuring the non-Hermioan observables, where the perturbations-induced imaginary part due to the measurement apparatus can be ignored and thus the preselection on measurement basis is needless. In other word, the uncertainty and randomness are not due to the selections by measurement apparatus, which mean the Green function here does not contains a self-energy term. In many-body microscopic case, the projective measurement suppress the long-range entanglement but does not cause the selections on measured states due to the effect of unique Markovian bath. Another reason is that we choose a time-independent basis $\langle\rho_{0}|\rho_{0}\rangle$. Consequently, the mapping $|\rho_{0}\rangle\rightarrow|\rho^{\prime}_{0}\rangle$ does not cause any new conservations. This indeed reflects the existence of response to the global quench in many-body microscopic case, just like the complete thermal state in late stage of single-body quasiparticle case, i.e., the common characteristic between the Markovian system in single-body quasiparticle case and the unique Markovian bath in many-body microscopic case. These are proved by the expectation (Green function) and variance of non-Hermitia observable[180] as well as the Jensen-Shannon divergence (JSD) in this article.
II Effect of measurement in bipartite configuration
Lindblad master-equation[16] for dissipative Markovian process, consisting of a coherent interacting term and a local dissipation term, and where measurement outcomes are averaged out, is most appropriate[13, 15, 14, 17] in the absence of detection of quantum jumps, in which case the system jumps to the state $|\psi(t+dt)\rangle=L\sqrt{\gamma dt}|\psi(t)\rangle$ with a probability of $\gamma dt\langle\psi|L^{{\dagger}}L|\psi\rangle$ ($L$ is the jump operator and $\gamma$ is the measurement rate). In this case with the presence of non-Hermitian skin effect, we can suspect the unique external disorder help the system to retain initial information, in opposite with the quasi-particle effect. This corresponds to the densities of the class of many-body microscopic ($mm$) $\rho_{mm;\{i\}}$, which is a Hermitian matrix and characterize the robustness of initial state revival as well as the bulk disentangling liquid which prohibits the diffusion of quantum mutual information (correlation) among the whole system globally. The counterpart of $\rho_{mm;\{i\}}$ is defined as $\rho_{sq;\{i\}}$, which is a non-Hermitian matrix. Here the subscript $sq$ referes to single-body quasiparticle. For $\rho_{sq}$, the non-Hermitian skin effect is absent, and there is delocalization as well as thermalization due to the non-Hermicity induced by $\mathbb{Z}_{2}$-type field, before the thermal equilibrium where the system becomes exactly Hermitian and local with respect to the unique environmental disorder (thermal bath) which should be distingushed from the above-mentioned external disorder for $\rho_{mm}$ Also, for $\rho_{sq}$, there is a diffusion (nonreciprocal dissipation) of quantum mutual information (carried by heavier carrier) among the whole system globally, and thus there will not be disentangling liquid within the bulk but only be a single disentangling boundary. Consequently, such correlation is system-size-dependent, or, extensive. This meets the label of single-body quasi-particle whose lifetime entirely controlled by the unique environmental disorder, and it is impossible for the multi-bound-states to appear like the case of $\rho_{mm}$. Importantly, even the non-Hermitian density $\rho_{sq}$ is diagonalizable due to the existence of non-defective degeneracy (NDD). While for the non-Hermitian operators that contains only the defective degeneracy, it becomes nondiagonalizable at certain momentum termed exceptional point. This means the set of eigenvectors that support non-defective degenerate span a full Hilbert space, in contrast to the defective degenerated one. Importantly, the non-defective degeneracy of $\rho_{H}$ preclude the symmetry-protected many-body localization and guarantees the global symmetry (chaos), thus its complex eigenenergies is different from that of a quasiparticle in a system with topologically enclosed exceptional point, such that its imaginary part (inversed lifetime) is directly determined by the unique environmental disorder from thermal bath (unlike the regular quasiparticles induced from a quantum jump process), and we further found that its real part of eigenenergies correlated (uncorrelated) with the imaginary part outside (inside) the NDD block. While for regular quasiparticles the imaginary part of eigenenergies close and away from the exceptional points are always related to different sources. It is found that the topologically stablity of exceptional points requires two or higher dimensions for a nontrivial topological index and can only be created or annihilated in pairs[82], thus the defectively-degenerated system containing exceptional points requires additional dimension compares to the non-defectively degenerated one. This is in consistent with the NDD eigenvectors constructed in our configuration, where the forward-scattering-like[172] eigenvectors in degenerated region preventing the formation of many-body localization, and the superpositions of NDD eigenvectors form a null-(sub)space whose correlation inside is extensive and thus guaranteeing the global symmetry. Such extensive correlation, or the dependence on the size of global system, is in analogy with the topological invariant with noncontractible path in momentum space[82], unlike the contractible path for a non-Hermitian system with defective degeneracy (exceptional points enclosed by the path).
In thermodynamic limit with projective measurement, the correlations as well as entanglement entropyis bounded (unbound) for ergodic (coherent) system, as detailed below in $sq$($mm$)-case.
II.1 Long-range entanglement and thermal equilibrium steady state
The trajectory of discontinuous conditioned stochastic evolution with quantum jumps becomes the continuous diffusive quantum trajectory in the limit of strong dissipation with an infinitesimal Wiener increment (the Markov limit which invalidate the photon detection by white noise). The Wiener increment has a zero mean over the evolution described by stochastic Schrödinger equation instead of Lindblad master equation. Such a stochastic continuous measurement (which initially acts locally on the quantity like occupation number or spontaneous emissions in an optical lattice) mainly contributed by the Wiener process (Brownian motion) with the transition from jump evolution (stochastic) to diffusive evolution (Ito stochastic)[45, 91, 34, 41, 51, 50, 90]. This also accompanied by a transition from ballistic propagation (spreading) of quasiparticle (entanglement) to diffusive spreading of, e.g., energy. Despite initially the discontinuous local measurements on the conditioned evolution of jumps (i.e., low emissio rate in which case the Born-Markovian approximate is valid) will hinder the thermalization (decoherence), the raised emission (dissipation) rate or strongly diffusive transport through the collision-induced quasiparticles[50, 51] (or other nonlocal carriers), guarantee the thermalization in final stage[28, 91, 34]. The ballistic (and uncorrelated) spreading of informations is suppressed by the local measurements as well as the unitary evolution[45, 46]. The continuous homodyne or heterodyne non-projective measurements are of such detection[34, 41, 91]. In some cases, the white noise limit leads to maximally mixed state with maximal entanglement (e.g., infinite two-point correlation[43, 58]). While the local projective measurements will suppress the unitary-evolution-induced short-range entanglement, the local projective measurements with unitary evolution can contribute to the generation of long-range entanglement or a global (topological) symmetry[44] at low-temperature, and being area-law scaling for steady state. In the other hand, a mode involving mutually commuting particles (repeated action) among a long-range[39, 51] that storing the same informations, generally has long-range entanglement, like the nonlocal condensate with global phase locking[51], and nonlocal measurement is applied for such information. In fact, long-range (short-range) entanglement usually emerges in local (thermal) state with area (volume) law scaling entanglement[81, 44, 39], like the Bose-Einstein condensate dark pure state (non-Abelian anyons). Note that here the nonlocal condensate is out-of-thermal-phase superposition coherence. Such mode applied in Born-Markovian approximation implies the absence of bound state or coherence and the presence of weak-dissipation-coupling-induced spontaneous emission and collision. In the absence of spontaneous emission the entanglement is generated by the global incoherence which is exactly of the non-Hermitian picture depicted in the early stage of $sq$-case; An circuit realization is achieved through the global incoherence noise which acts uniformly on each every qubit[55]. While the above-mentioned decoherence case is depicted in the late-stage of $sq$-case.
While approaching the continuous-limit leads to measurement-induced entanglement transition in single-particle picture with area-law saturating entanglement and maximally mixed steady state with maximal measurement (or dissipation) rate upto thermodynamic limit, a volume-to-area transition is visible at finite time for the discontinuous measurement (without unitary conditional feedback), termed as pseudo-skin effect under open boundary condition[56, 52]. Such pseudo-skin effect as well as dynamical entanglement transition is manipulated by the time step of quantum jump, instead of the disorder, conditional feedback or measurement.
Except the measurement-induced entanglement transition, there is another kind of supression on long-range entanglement, which is by the conditional feedback (locally operates on the unpostselected trajectories) with the trajectory-average consistent with the Lindblad master-equation approach[12, 13, 28]. Such feedback-induced skin effect can happen in both the noninteracting (free fermion) system and chaotic interacting system, and requires continuous measurements (white noise limit), thus results in unconditioned diffusive trajectories with short-range area-law entanglement and suppress the measurement-induced entanglement transition (i.e., from diffusive logarithmical scaling entanglement to saturating one)[52, 39].
Similar to the skin effect of non-Hermitian system feedback-induced skin effect results in area-law entanglement for the steady states concentrate at edge, and there may even be multiple edges for unitary dynamics. Differently, in the absence of spin effect ($sq$-case), each subsystem thermalizes following the volume law scaling before approaching the steady state which is extensive with subsystem size. Thus these are two distinct kinds of disentangling.
The repeated local measurement with local feedback cause the averaging over the coherently superposed trajectories (nonindividual). Specifically, at generic unital quantum channel the identity operator becomes stationary with short-range entanglement. For such (passive) measurement with unconditioned dynamics, there is an average on the trajectory-ensemble level, and essentially, this average is a coherent superposition instead of a statistical mixture since the latter equivalent to a ensemble of identical copies (individual trajectories) each in a pure state. Comparing the Lindbladian master equation with the non-Hermitian Hamiltonian, the conditional feedback only operates on the trajectories deviated from postselection, i.e., the quantum jump processes (collective dissipation with each individual quantum trajectory subject to stochastic loss events), $\sum_{i}\gamma_{i}L_{i}\rho L_{i}^{{\dagger}}$ with $\gamma_{i}$ and $L_{i}$ the dissipation rate and jump measurement operator, respectively. The quantum channel map to itself when $[\rho,(\sqrt{\gamma_{i}}L_{i})^{{\dagger}}]=0$, such that $\sum_{i}(\sqrt{\gamma_{i}}L_{i})\rho(\sqrt{\gamma_{i}}L_{i})^{{\dagger}}=\rho$. While system temperature can be affected by the measurements only when $[H,(\sqrt{\gamma_{i}}L_{i})^{{\dagger}}]$ is nonzero[80, 87].
II.2 Short-range entanglement and nonequilibrium steady state
For $mm$-case the short-range entanglements (with nonlocal measurement) is necessary where there are coherences with weak dissipation (i.e., the non-Markovian dynamics with robust memory effect) and finite-time thermalization (in terms of energy spectrum, this can be evidenced by the gapped structure and variant number of bound states[47, 10]). Thus different to the above cases of density-bath (decoherenced and disentangled), the strong system-bath coupling produce bound states with high fidelity. Another phenomenon in the limit of closed system is tunneling as a stochastic process where the non-perturbational resonance at some certain local spatial sites is overwhelming compares to others due to the statistical fluctuations of the on-site disorder. This give rise to the non-Hermiticity through the mixing of non-adjacent eigenstates and all-to-all coupling as can be evidenced by the complex eigenenergies[13, 56, 173].
Note that the all-to-all limit (Dicke limit) corresponds to the lowest effective dimension (where, e.g., the long-range hopping prompt the homogenization in density distribution), or infinitly large system size, where the open boundary condition and periodic boundary condition produce similar dynamics, similar to the case of strong-disorder-induced dynamically stable non-Hermitian many-body localization with the evolution dominated by area-law scaling. Also, in all-to-all limit, the localization length as well as the density imbalance of the particle distribution becomes nonextensive. Under periodic boundary condition, there is uniform density distribution in late-time steady state[56] (without imcompressible density or domain structure in spatial space). While in the presence of skin effect, which is compatible with many-body nature as long as it is not in the all-to-all limit and involves finite interacting effect (dissipation), i.e., in the nonreciprocity limit of Hatano-Nelson model with biased (imbalanced) density distribution, the entanglement phase transition only exists for weak disorder under open boundary condition before it reaches the late-time steady state.
Through Loschmidt echo, a robust initial siate fidelity reflecting such high-overlap-induced large amplitude (be independence of bath). This should be distincted from the amplified local oscillator in coherence field in homodyne limit with continuous measurement, where the coherence is tightly related to the bath (infinite photodetection rate) and obviously not a closed system. The complex coherence pattern as well as the emergent bound states can be visualized by the Loschmidt echo, as detailed below. Additionally, the conserved quantities (like the bound state) exhibit linearly (in Hilbert space dimension) large lifttime (relaxation time) and thus logarithmically slow the transport[48] and lower the correlation[43], deviate away from the white-noise (Markov) limit.
III Trace distance and Majorization
The off-diagonal fluctuation, which plays a subleading role in ETH) vanishes in long-time limit, and thus the off-diagonal elements exponentially small in system size. A completely thermalized final state has zero off-diagonal part, and can be described by a gapless phase[10, 19] As a result, the deviation of the density from the one in diagonal ensemble, $\rho(t)-\rho_{d}=|\psi(t)\rangle\langle\psi(t)|-\sum_{i}|\langle i|\psi_{0}\rangle|^{2}|i\rangle\langle i|$, satisfies
$(\rho(t)-\rho_{d})^{2}=\sum_{i\neq j}\langle\psi_{0}|i\rangle\langle j|\psi_{0}\rangle e^{i(E_{i}-E_{j})t}|j\rangle\langle i|\sum_{k\neq l}\langle\psi_{0}|k\rangle\langle l|\psi_{0}\rangle e^{i(E_{k}-E_{l})t}|l\rangle\langle k|,$ (1)
whose long-time average reads
$\langle(\rho(t)-\rho_{d})^{2}\rangle_{t}=\sum_{i\neq j}\langle\psi_{0}|i\rangle\langle j|\psi_{0}\rangle|j\rangle\langle i|\langle\psi_{0}|j\rangle\langle i|\psi_{0}\rangle|i\rangle\langle j|$ (2)
$=\sum_{i\neq j}|\langle\psi_{0}|i\rangle|^{2}|\langle j|\psi_{0}\rangle|^{2}|j\rangle\langle i|i\rangle\langle j|=h^{2}-\sum_{i}|\langle\psi_{0}|i\rangle|^{4}(|i\rangle\langle i|)^{2},$
where $h=(\sum_{ij}|\langle\psi_{0}|i\rangle|^{2}|\langle j|\psi_{0}\rangle|^{2}|j\rangle\langle i|i\rangle\langle j|)^{1/2}$ is a Hermitian full-rank matrix. The variance of eigenvalues of $h$ is nearly three times of that of $h^{2}$. Note that $h\neq\rho_{nd}$ with $\rho_{nd}$ be a idempotent matrix $\rho_{nd}:=\sum_{ij}\langle i|\psi_{0}\rangle\langle\psi_{0}|j\rangle|i\rangle\langle j|$, but ${\rm Tr}h^{2}={\rm Tr}\rho_{nd}^{\gamma}\equiv 1$. The last term in above relation satisfies $\sum_{i}|\langle\psi_{0}|i\rangle|^{4}(|i\rangle\langle i|)^{2}=\rho_{d}^{2}$, where the coefficient ${\rm Tr}(\rho_{d}^{2})=\sum_{i}|\langle\psi_{0}|i\rangle|^{4}(<1=\sum_{i}|\langle\psi_{0}|i\rangle|^{2})$ defines the purity and tell the deviation of $\rho_{d}$ from a pure state. The corresponding traces satisfy
$\langle{\rm Tr}[(\rho(t)-\rho_{d})^{2}]\rangle_{t}={\rm Tr}[(\rho(t)-\rho_{d})^{2}]\ (\forall t)={\rm Tr}h^{2}-{\rm Tr}\sum_{i}|\langle\psi_{0}|i\rangle|^{4}(|i\rangle\langle i|)^{2}=1-\sum_{i}|\langle i|\psi_{0}\rangle|^{4}>0.$ (3)
While ${\rm Tr}(\rho(t)-\rho_{d})=0\ (\forall t)$, which corresponds to the vanishing fluctuation (off-diagonal part) for the system of $\rho_{H}$. For the $sq$-case, there without any atypical state (nonthermal and nonergodic) and the initial state revival (i.e., back to the nonequilibrium state) cannot be found, which corresponds to the case that the subsystems of the nonequilibrium state be a nearly vanishing fraction of the total states. This is a scenario much more closer to equilibrium compares to the one discussed in Ref.[19]. Considering a bipartition of the restriction-allowed states into the system part and environmental part, smaller the dimension of system part $d_{s}=(\sum_{i}|\langle i|\psi_{0}\rangle|^{4})^{-1}$ (inversed purity) is, shorter the distance between $\rho(t)$ and $\rho_{d}$. Reminiscent to the distance between a state of very small subsystem $\rho_{S}={\rm Tr}_{E}\rho_{{\rm pure}}$ and the canonical state $\rho_{{\rm can}}=\langle\rho_{S}\rangle={\rm Tr}_{E}\rho_{{\rm mix}}$. For a global restriction, each pure state of the system becomes close to the maximally mixed state $\rho_{{\rm can}}={\bf I}_{s}/d_{s}$[21, 22, 23]. The process of such proximity can be realized guadually by enlarging the kinematical restriction to the size as large as the universe. However, we found that such averaging feature can be realized without the exponentially small dimension of subsystem compares to the rest part, as long as there contains a NDD nullspace, i.e., the eigenvector set span a full Hilbert space.
While for the many-body microscopic system (the $mm$-case), each subspace, which is also the smallest subsystem, is spanned by a rank-1 pure state, then we found that each smallest subsystem is evolved to the equilibrium which can be seen from the stable pattern of fidelity (), and such equilibrium state does not thermalized with respect to the environmental disorder (thermal bath) and be able to coherent with other subsystems (in arbitary size). Thus final equilirium state behaves as a mixed state governed by non-Abelian symmetry, and this is in agree with the statement of Refs.[19, 30, 21]: For a sufficiently small subsystem (thus a small fraction of the total states) almost every pure state is locally indistinguishable from the mixed state with entangling (whose trace over the environment part is a thermal canonical state). In other word, the massive entanglement between the very small subsystem and its effective environment render the proximity from a pure state to maximally mixed state with maximal entanglement, and this is also related to the majorization of pure states[25], e.g., the arbitary state with probability distribution away from uniform (equilibrated) distribution majorize the maximally mixed state[24, 25] which has a largest entropy. We found that, this characteristic of small subsystem can be seem from the subsystems of the NDD regime, such that arbitary subspaces or subsystems of the NDD regime exhibit no coherence with the arbitary piece of the rest part of global system. This can be seem from the thermal character and the persistive initial-state revival (large recurrence). Importantly, in our configuration, it requires not exponentially large dimension of the environmental part compares to that of the subsystem. This can also be understood as a large effective dimension is hidden in a system containing the NDD.
A pure state can be entangled or unentangled, depending on it is separable or not. This correspond to the two kinds of subspaces (complex density matrices) of the mm-case or sq-case (in different bases), where the former are unentangled while the latter are entangled. In other word, there is a subjective lacking of information regarding the global system for the pure states of mm-case ccompares to the sq-case, The regain of such subjective information (weight on each subspace) is possible for small enough subsystem[21], or by move to the NDD nullspace (where the eigenstates become real and the corresponding subsystems therein become Hermitian). Thus, such subjective information is essential for the entanglement, and as a result, the entanglement is possible to be preserved even in Gibbs stationary states of some non-local Hamiltonians, whose quantum channel can be artificially designed[26]. In bipartite configuration, a bipartite pure state reads $|\psi\rangle=\sum_{i}^{D}\lambda_{i}|\psi_{S}\rangle\otimes|\psi_{E}\rangle$, and the reduced states are $\rho_{S(E)}={\rm Tr}_{E(S)}|\psi\rangle\langle\psi|=\sum_{i}^{D}|\lambda_{i}|^{2}|\psi_{S(E)}\rangle\langle\psi_{S(E)}|$. The pure bipartite state is maximally entangled, if the reduced density matrix on either system is maximally mixed, and the corresponding Schmidt coefficients be $\lambda_{i}=\sqrt{1/D}$ and the purity of each subsystem is $\sum_{i}|\lambda_{i}|^{4}=1/D$. In opposite case, the separable pure states has $|\lambda_{i}|=\delta_{i,1}$ the each reduced system as well as their separable product is pure state, such that $\rho_{S}\otimes\rho_{E}=|\psi\rangle\langle\psi|$. In sq-case, a entangled pure state $|\psi_{1}\rangle$ is possible to be transformed into another state entangled $|\psi_{2}\rangle$ as long as it satisfies ${\rm Tr}_{S}|\psi_{1}\rangle\langle\psi_{1}|\prec{\rm Tr}_{S}|\psi_{2}\rangle\langle\psi_{2}|$ (we denote the eigenvalues of these two matrices made up the vectors $|E_{1}\rangle$ and $|E_{2}\rangle$, respectively). This transformation can be implemented by the Local Operations assisted by Classical Communication (LOCC)[27, 24]: a single local measurement on the system part and a nonlocal unitary operation on the environmental part. Then it is guaranteed that $|E_{1}\rangle\prec|E_{2}\rangle$, i.e., the state $|E_{2}\rangle$ majorize the state $|\psi_{1}\rangle$ since the latter is disordered stronger than the former, (such that their entropies $S(|\psi_{2}\rangle)<S(|\psi_{1}\rangle)$).
For the densities of $mm$-case, for $i$-th subspace, $|E_{i}\rangle=(0,0,\cdots,0,1)^{T}$. $|E_{i,j,k,\cdots}\rangle\prec(0,0,\cdots,0,1)^{T}$ where $i,j,k,\cdots$ denote the distinct subspaces, and this majorization relation is valid as long as it include more than one subspace. Specfically, there is super-majorization for $|E_{i,j,k,\cdots}\rangle$ including all the subspaces, such that $\sum^{D-1}_{i}E_{i}(\rho_{mm;c_{1}})<\sum^{D}_{i}E_{i}(\rho_{mm;c_{1}})=1$. This is consistent with the unentangled subspaces of the $mm$-case, as the coherence between the arbitary subsystems outside the NDD regime produce the non-local symmetries (bound state in infinite amount) where the induced short-range entanglement overwhelming the local measurement-induced disentangling (volume law to area law transition; which appears in the $sq$-case).
IV Entanglement and measurement
IV.1 $sq$-case
The $sq$-case corresponds to the monitored system with random quantum jumps (discontinuous), where the measurement results are unknown due to the average in trajectory ensemble. Also there are long-range entanglements with adaptive dynamics in this case, before the averged state is being completely thermalized (dephased), with respect to measurement basis, into the Hermitian local states.
Thus the entanglement transitions from volume law to area law (mainly refer to the one induced by the skin effect) is more significant for the conditioned trajectories, since the entanglement is contingent upon the measurement outcomes[36] and diminished once the measurement record lost.
Considering the trajectory-averaged entanglement, the reduced density matrix on $i$-th eigenvalue $R_{i}=\begin{pmatrix}\lambda_{i}&\nu\\ \nu^{*}&1-\lambda_{i}\end{pmatrix}$ always has the eigenvalues majorize its diagonal elements as long as the off-diagnal element is nonzero ($\nu\neq 0$): $(\frac{1}{2}(1+\sqrt{(1-2\lambda_{i})^{2}+4\nu\nu^{*}},\frac{1}{2}(1-\sqrt{(1-2\lambda_{i})^{2}+4\nu\nu^{*}})\succ(\lambda_{i},1-\lambda_{i})$. The classical Shannon entropy provides a upper bound on the entanglement entropy, $-\lambda_{i}{\rm ln}\lambda_{i}-(1-\lambda_{i}){\rm ln}(1-\lambda_{i})\geq-{\rm Tr}[R_{i}{\rm ln}R_{i}]$. Note that for Shannon entropy, the eigenvalues always be normalized to be $\sum_{i}\lambda_{i}=1$. This entropy uniquely determined by $\lambda_{i}$ is also called Boltzmann one-body entropy[35], which refering to the single-particle Hilbert space. As also reflected by the Loschmidt echo, the free-particle in single-particle picture exbihit much weaker ability in quantum information memory, due to the lacking of interaction (with reservoirs).
The $sq$-case and $mm$-case reflect the entanglement (sub)additivity and symmetry additivity, respectively. As a result, the former has a much larger measurement rate, and exhibits more pronounced randomness.
The disentangling effect induced by Abelian symmetries, like $Z_{2}$ gauge symmetry or $U(1)$ symmetry, appear only in the $sq$-case after long time evolution. Although such Abelian symmetrie is fragile to the dissipation processes[37], fluctuation of order parameter, charge diffusion[38], or noisy-driven biased random walk, the symmetry breaking is avoided by the effectively low-dimensional structure of the eigenspace where the eigenstate coalescence (as happen in delocalized phases) is absent. In other word, the non-defective degenerate structure in $sq$-case effectively avoid the dynamical instability of the delocalized eigenstates (in the early stage of evolution), and consequently there is no emergent symmetry sectors and entanglement between each subspace is robust. As the non-Hermiticity dominate in the early stage of (non-unitary) evolution, the nonlocal measurements on the individual conditioned trajectories can generate entanglement[36, 39], but different to the kind of adaptive dynamics as mentioned above (which is of the Hermitian case).
While before long-time limit, there is global quantum correlation (i.e., global entangling) throughout the whole system, where there is continuous non-Abelian symmetries, like the SU(2) symmetry, and thus allowing the long-range entanglement and logarithmically large entanglement. Typically, this is the case for closed quantum system, where the quantum mutual information diffuse throughout the whole system during evolution, with extensive correlation (or entanglement), and finally realize thermalization with the vanishing off-diagonal fluctuation and non-Hermiticity. Opposite to this, the $mm$-case corresponds to the discrete non-Abelian symmetry, where there is (relatively) short-range entanglement with local unitaries. As a typical example, for the quantum jump term as mentioned above which describes the measurement with an unknown result (to be averaged out), it turns to be a strongly symmetric channel if the jump operator can be decomposed into a form that is diagonal in the space of each irreducible representation (irrep) and not on the multiplicity space[30]. Thus it allows both the cases of non-Abelian symmetry and Abelian symmetry, which has more than one and only one state in each irrep, respectively.
In the presence of large dissipation and decoherence arised by the (weak) system-bath interaction, the entanglement growth is suppressed such that the dephased states have area law scaling due to the ergodicity, and the single-particle-picture is well preserved by low-effective-dimensionality of the eigenspace as mentioned ablve. For an arbitary subsystem which must behave as a open quantum system, the rest part can be treated as a thermal (or Gaussian) environments which can be expressed as a product of all the (commuting but long-range entangled) thermal states therein. Due to the definite eigenvalue for each subspace, the operator expectation, which is a linear quantity (linear on the time-dependent stochastic states), can be retrieved from a master-equation[39, 40]. One of the monitoring protocol for weak measurement is the phase locking between adjacent sites[39] which is consistent with the low-effective-dimension structure, and low-effective-dimensionality throughout the whole system. Similar to the special structure of the NDD eigenvectors, the phase locking preventing the many-body localization by breaking the local U(1) symmetry. Note that the phase locking here is of the Markov process accomponied by the spontaneous emission (quantum jump). Consistent with the fact that the phase locking is closely related to the long-range entanglement and Born-Markovian approximate (which is applicable in the case of weak system-bath coupling and without bound state[47, 39]), there is a strong dependence on the subsystem size for the entanglement entropy growth, which move to the area law scaling in the regime where the dephasing dominates (with strong system-bath coupling).
In $sq$-case, the thermalization is allowed by the diffusive spreading of initial state information (or correlations) over the entire system (upto long-time limit) and meanwhile small local subsystems are possible to loss the memory about initial information and relax to thermal state under weak disorder. This corresponds to the non-local measurement, e.g., the measurement ideally affects all the bonds in a spin chain model, which will leads to superballistic spreading (of entanglement or quantum information) under nonunitary evolution[46]. Such superballistic spreading has a faster-than-linear entanglement growth, which appears in the initial stage of spreading[50] where there is equivalent bipartition with maximal thermodynamic entropy. While in the presence of nonlocal measurement and unitary evolution, there is ballistic spreading with linear entanglement growth following the volume law. Distinct to the former, the ballistic spreading involving the process with information spread from one heavy carrier to another heavy carrier through the carrier-carrier interaction, such that the entanglement is generated between those carriers (with information scrambling). An example of such carrier (a local degree of freedom) satisfying the volume scaling law is the spin[50, 15]. In contrast to nonlocal measurement, the local measurement will not prohibit the global thermalization, whose impact is exponentially localized[49]. Importantly, the non-local measurement cause the ballistic entanglement spreading (and diffusive entanglement spreading throughout the system in the late-time stage [15, 50]) in $sq$-case where the perturbation from local projective measurement is avoided by the effective single-particle picture, coexist with the diffusive energy (or information carriers) transport throughout the whole system. Except the continuous level distribution[92], we further found that the equally-spaced eigenvalues (of the thermal states) also reveals the invalidity of perturbation theory. In this case, the only local observable is the total energy, thus there is no ballistic transport in terms of ballistically traveling quasiparticles[50], like photon is absent, and the only way for energy transport is diffusion (through strong scattering).
IV.2 $mm$-case
The $mm$-case corresponds to a constrained Hilbert space with extensive entanglement, where the coherence pattern prohibits the diffusive spreading of correlations over the entire system and thus exhibit robust initial state fidelity (and purity of the densities). Note that here the constraint on Hilbert space, which results in fail of global thermalization and the initial fidelity, should of the same type of the repeated local measurement (on an individual bond or qubit), and the constrain on Hilbert space as well as the reduction on entanglement depends on the frequency of local projective measurement. The complexity of coherence field is enhanced with the raising frequency of local measurement, and the global thermalization (with saturated entanglement growth following area law) cannot achieved until at the maximal frequency which involve the off-diagonal densities such that the system turns to, again, a local tensor product structure. Thus unlike the above $sq$-case, the coherence field (local oscillation) in $mm$-case generates the infinite number of bound states (quasiparticles) that allows the ballistical transport at finite energy. But the merged Loschmidt echo indicating the zero-energy limit with saturating scaling instead of ballistical one. In fact, the static entanglement entropy (like the Renyi entropy) soly depends on the variance of boundaries of Loschmidt echo, such that each subspace of $mm$-case has zero entropy which behave exactly like a pure state, and for arbitary subsystems, the Renyi entropy (up to arbitary order) of $sq$-case always higher than than of $mm$-case if we donot consider the dynamics of entanglement (i.e., without consider the increasing complexity during unitay evolution).
In the $mm$-case, a large amount of bound states (mode) emerge discretely as can be seen from the Loschmidt echo, where the superpositional coherence is allowed between the open subsystems. Also, the discretely emergent non-Hermitian skin effect with unthermalized-type disentangling cause the absence of measurement-induced entanglement transition (volume-law to area-law)[28]. The weak disorder and the non-Hermiticity result in the delocalization in $mm$-case, which means the system is free from the constraints of integrability and many-body localization, and, in the mean time, without thermalization and global quantum correlation or entanglement[15].
Arbitary densities $\rho_{mm;\{i\}}$ therein (with ${\rm Num}\{i\}>1$ and thus be the mixed state) behave as an ensemble of pure states[31], which act as an open subsystem interacting with infinite reservoirs[31], and thus the coherence plays the dominating role (with the off-diagonal part lacking large fluctuation). The skin effect is available for arbitary strength of non-Hermiticity, thus it produces the disentangling (suppresses the entanglement and thermalization) in a different way compares to the measurement: The non-Hermiticity is realized by the constriction to a subspace in finite size (symmetry sector) with the conditioned (postselected) dynamic of quantum trajectories on the measurement outcomes, where exponentially many measurements are required. The measurements on the conditoned dynamics in non-Hermitian system return null results due to the lacking of average over trajectory ensemble as well as the absence of randomness in the measured dynamics.
In the presence of finite dissipation, while the non-Hermiticity refers to the microscopical continuous measurements with the conditioned evolution without quantum jump, there are discontinuous conditioned evolution with quantum jumps for the whole picture of $mm$-case. Those involved quantum jumps can be directly evidenced by the repeatly merged Loschmidt echo, which also refers to the coherent field (local oscillator) in a system like optical cavity. $sq$-case always process a linearly-in-time (volume law) to saturating (area law) transition, $mm$-case (without postselection) process a volume-law entanglement without saturating[135] and exhibit persistent fluctuations with the increasing complexity of coherence configuration, i.e., the set of local Hamiltonian (disorder-free). It is such locality and coherence that guarantee the fidelity (as well as the subsystem purity).
IV.3 Realizations under measurement
The above-mentioned dissipation with measurements, which allows measurement-induced entanglement transition as well as other continuous phase transitions, always prompt the decoherence due to the rised system-bath coupling in frequency, consistent with the behavior of a Markovian open quantum system. In other word, there is a single local operator projected into the whole system which is Gaussian localized (ergodic and has extensive entropy). For a Gaussian distribution implied by the central limit theorem, the randomness and mutual independence is hypothesis[59] instead of a rigorously proven fact. One reason is the exponentially localized perturbations from local measurement will not affect the global locality and thermalization. This should be distinguished from the nonreciprocal dissipation (the origin of skin effect), which give rise to discontinuous phase transition.
The magnitude of imaginary eigenenergies directly determines the instability of the delocalized phase against the absorption or emission of modes (quantum jump)[13]. Larger the imaginary part, smaller the uniform damping[138]. This is also in against with the no-jump condition for non-Hermitian dynamic in continuously measured quantum system, and thus effectively suppress the entanglement through the local measurement. Further, the weak and frequent (strong and sparse) measurement results in the long-lived pure state (mixed state) with suppressed (enhanced) entanglement following the area-law (volume law). The former under weak measurement (Markovian proces) occurs with broken parity-time ($\mathcal{P}\mathcal{T}$) symmetry [61, 60]; While the latter under strong measurement requires certain symmetries, like the time reversal symmetry[13], which is also consistent with the suppressed imaginary part of eigenenergies by localization (on-site disorder). The strong measurement usually refers to the projective one, during which process the system-probe coupling is much stronger than the energy scale of system, and consequently a weighted trajectory process volume-law (steady-state) entanglement. For the former case with weak measurement, the global localization caused by the dense measurement is the so-called quantum Zeno effect[108, 66, 138, 112, 136]. The Zeno effect, unique to the $sq$-case, is accompanied by the continuous measurements and continuous strong dissipations which suppress the fluctuation as well as heating up of the system[112, 63, 134, 135, 148] as a property unique to the nonequilibrium engineered dynamics, and thus protecting the topological order therein[113, 81], and the average over environmental non-Markovian noise realizations cause the non-Gaussian correlation under nonperturbative treatment. Note that here the disentangling (area law) due to the Zeno effect is distinct from that due to the projective measurements in $mm$-case, as the former stabilizes the measurement-induced phase transition while the latter hinder it. Here the suppression on fluctuation implies the common effect between the continuous strong dissipations and the on-site disorder which is capable to leads to localization even in non-Hermitian many-body system[40]. The diffusion-induced continuous dissipations here is nonlocal and related to the widely spreading quantum correlations. Another property similar to the Zeno effect is the nonstabilizerness whose additivity guarantees its global nature and can even be produced by a unitary operator by averaging over the ensemble of pure state trajectories of a full unitary group[110, 111]. The Zeno effect is most significant in the later stage of $sq$-case, whereas in early stage the large system size can enhance the robustness of entanglement grow against the classical stochastic dynamics by the quantum jumps[136, 153], i.e., lower the effect of dissipation. Note that the heating loss (suppression on entanglement production) here is different to that in the $mm$-case, but relys on the robust initial state property with non-Hermitian collapses and revivals[139], similar to the quantum cooling procotol of Ref.[146] which is by reducing the excited states back to ground states. While the Loschmidt echo directly indicates that the discrete symmetries only exists in $mm$-case with the jump-induced local dissipations (accompanied by the local dephasing and oscillatory behavior[112]) and dissipative coupling via the ballistic transports and synchronization, which is unique to a many-body system and representing the robust initial-state fidelity under the measurements at a finite rate[88, 66, 109, 151]. Except the many-body nature, the infinitely long time needed to loss the initial-state information also consistent with the thermodynamic limit in $mm$-case, where there are thermal baths holding well-defined temperature.
The entanglement transition arises from the sparse-to-dense measurement is invisible in the presence of skin effect or density-bath coupling through the measurements which results in trajectory-averaged evolution. In other word, it requires some temporal randomness[60, 65] to realize measurement-induced entanglement transition, e.g., in ergodic and nonintegrable system. Such randomness can be caused by the interacting effect (such that the free fermions system under weak measurement does not follows volume-law scaling[28, 136]), including the spin-orbit coupling between the up-spin fermions and down-spin fermions[61]. Here, the average over (trace over) trajectories (measurement outcomes) results in the dynamics of mixed state density matrix described by Lindblad master equation. Such statistic mixture necessarily produce area law short-range entanglement in the open boundary system, and allows: 1) The temporal randomness in dynamics or measurement outcomes under weak measurement, which is important in keeping mixed state and developing volume-law entanglement; 2) Unitary evolution; 3) Applications of the quantum channel[60, 81, 78, 77]. Despite the quantum information can be well preserved by a local quantum channel through coherent measurement[77], or by the exceptional points (defective degeneracies), there are nonlocal quantum channels in terms of ensemble average where the combined effect of non-Hermiticity and many-body property cause the weak postselection (where the nonlocality dominates and preclude the exceptional points), and the Lindblad master equation is the continuous limit of such a quantum channel[75, 60]. Note that such weak postselection case with significant randomness corresponds to the $mm$-case.
Distinct from the measurement-induced phase transitions deeply affect by the disorder in open quantum system coupled to a Markovian environment (i.e., the process of repeated weak nonlocal measurements)[13, 61, 60], the skin effect does not affected by the disorders. Importantly, in the presence of unitary conditional feedback together with the projective measurement, the system no necessarily prohibit the measurement-induced entanglement transition[62, 63], this generally corresponds to the case of weak measurement. Consistent with our above analysis, such weak measurement scheme, which is free from the skin effect as well as the induced disentangling, corresponds to the $sq$-case: The quantum state diffuse over the whole system and evolve coherently (Heisenberg evolution) under measurement, $|\psi\rangle\rightarrow e^{-iH\Delta t}|\psi\rangle$ with the time step $\Delta t$ (strength of system-probe coupling) be comparable with the time-scale (energy-scale) of the system. Consistent with the behaviors unique to the $sq$-case, the system and the measuring setups are collectively in a pure state (closed) while the system by itself is in a mixed state[60, 61]. Here the coherent evolution should be distinguished from the coherences in the $mm$-case which relys on skin effect and individual quantum trajectory. In $mm$-case the system by itself relax to a pure state in terms of coherence pattern,
As detailed in the discussion on $mm$-case, we consider the short-range entanglement as the main source of volume-law scaling before the thermalization (Hermitian limit). Refs[66, 67, 68, 69, 71] also suggest that the entanglement is suppressed by measurements upto a certain frequency for finite probability of measurement outcome. The projective local meaurement[91, 34, 28, 39] projects the cavity mode into photon-number eigenstates, with the resulting photon number operator be Hermitian and commutes with the Hamiltonian (which governs the unitary evolution of the stochastic states). Except the measurement outcomes, the stochasticity could originates from the random space-time sites of noises (i.e., the quenched disorder which will leads to localization and entanglement suppression), and such stochasticity disappears for Poissonian measurement which take place randomly at each space-time site regardless of the specific probability[66]. While when the probability takes effect, the space-time sites in unitary circuit are biased by the nongeneralized measurements, where both the measurements on sites (through gates) and the measurement outcomes are specified (e.g., by postselection), and results in nonunitary dynamics[66, 60]. This is exactly the $sq$-case, which corresponds to a generic (nonnintegrable) system lacking the noisy and subject to global thermalization with continuous measurements.
Importantly, we stress that the full dynamics of $sq$ ($mm$)-case is nonunitary (unitary) in perspective of the trajectories and the entanglement growth, corresponding to the open system coupled to the nonthermal (thermal) non-Markovian (Markovian) environment with continuous (discontinuous) measurements, diffusion-(jumps-)induced nonlocal (local) dissipations, and weak (strong) system-bath coupling due to the incomparable (comparable) timescales, as described by the single-particle (many-body) picture. In other word, for $sq$-case ($mm$-case) under weak (strong) measurement which disallows (allows) the decoupling of sectors, there is the finite (infinite) temperature average or non-Gaussian (Gaussian) correlation that is nonlinear (linear) on state through definite (indefinite) heating up during evolution, together with an ill-defined (well-defined) bath temperature for the nonnintegrable (integrable) many-body system which leads to ergodicity or generic feature (nonergodicity or nongeneric feature) and the thermalization happened at a single trajectory level (trajectory ensemble level). It is also more easy to realize thermal equilibrium (nonequilibrium) in $mm$-($sq$-)case as evidenced by the short-range order (long-range order) and unitary (nonunitary) dynamics therein[136, 153, 141, 132].
In perspective of the generalized Gibbs ensemble and weak ETH in thermodynamic limit (due to the small fraction of unbiased and unpostselected nonthermal states), for $mm$-case, the countable quantum jumps can be treated as the weak perturbations for the breaking of integrability, localizations, and quantum coherence. In disordered phase of the $mm$-case, there are extensive localization length with large dissipation rate (thus, projective measurement in high frequency) and short-range hopping (thus, short-range entanglement). In $mm$-case, the strong Markovian (white) noise from environment will enlarge the time of relaxation toward the (unitary) Hermitian limit, exceeds the waiting time between quantum jumps, such that the averaging over noise trajectories allows the mapping onto Lindblad master equation where the jump operators appear in the incoherence (dissipative) part. It is such long timescale that bound the volume-law (ballistic) growth of entanglement. Such perturbation is too weak to causes thermalization and meanwhile the perturbation theory is validate by the significant (nonuniform) gaps between eigenvalues in this case (i.e., discrete levels due to the discontinuous jumps). Desipte the short-range entanglement is suppressed by the skin effect and noise which has the same effect with perturbations (jumps), the deterministic Heisenberg coupling and the kinetical constraints on Hilbert space due to the coherences results in local operators and the ballistic entanglement growth (as a result of interaction effect). Consistent with our finding, the diffusive fast relaxation is possible even in $mm$-case in strong-noise limit under weak-perturbative treatment[96, 97, 99, 100, 101], despite the diffusion here is suppressed by the skin effect. Note that the diffusion constant is inversely proportional to the relaxation timescale, noise amplitude (noise strength) or measurement (dissipation) rate, and disorder strength in strong (Markovian) noise case and weak (non-Markovian) noise case, respectively, where the squared disorder strength can be estimated as the second moment of noise in the latter case[103, 100, 96, 120].
While the integrability is completely broken by the infinite photon flux in $sq$ case, where the weak but frequently repeated measurement (also known as generalized or non-projective measurement), each as well as every qubit (uncountable infinite photon flux) being measured at each time step no matter how small the time step is, and allowing the entanglement transition[66, 70, 80]. Also, the Poissonian number statistics of the coherent local oscillator with infinitely large amplitude representing the photodetection or quantum jump counting is very large, such that it can be approximately as Gaussian (robust to interactions), by treating the field as a whole object and validating diverging two-point (auto)correlation function[43, 139, 136, 112], with local real-valued Gaussian noise (Wiener) increment and emergent quantum noise from the stochastic noise.
Thus there could be transitions between nonunitary and unitary dynamics: As mentioned in Sec.IV.1, a nonunitary dynamics with nonuniform continuous measurement can be guaranteed by a relaxation time shorter than that the waiting time between quantum jumps (or projective measurements)[90, 80], and the absence of interaction effect therein allows the measurement-induced phase transition as well as the criticality phenomenon[90, 118]. The high measurement rate (frequency) cause a short Zeno timescale and coarse-grained dynamics due to the incomparable timescales between system and bath [63, 121, 80, 90, 117, 96, 144, 137]. The fast relaxation to equilibrium corresponds to the small dissipation rate $\gamma_{i}$, low occupation fraction, and small noise (two-point) correlation in non-Markovian limit[96] which consistent with the weak noise limit and the Lindblad master equation dominated by the unperturbed (dissipative and uncoherented) part. Fast relaxation gives rise to the dynamical stability due to the suppressed imaginary part of effective non-Hermitian Hamiltonian, as well as the transition to localized phase at steady state in the long-time limit with uniform projective measurement (and thus becomes less sensitive to the measurement unlike the conditioned or postselected dynamics), where the equilibrium state (ensemble average of trajectories[90] or coherent superpositions of symmetry-breaking states[94]), becomes indistinguishable from the time averaged state[80, 93]. Thus, in conclusion, the $sq$-case involving two stages in evolution: In early stage there are weak but frequent measurements (continuously) and strong system-probe coupling[80], dominated by nonunitary (non-Hermitian) stochastic dynamics and algebraic decaying with gapless nature toward the pure state. In late stage where the system approaches the equilibrium steady state, it becomes more significant for the nonextensive character of the localization length and density imbalance of particle distribution due to the vanishingly small dissipation rate and system-probe coupling, in which case the (group) velocity is well-defined according to the long-range power-law hopping of free fermions[56] due to the absence of many-body feature, and meanwhile it is possible to realize unitary and Hermitian dynamics due to the emergent topological invariants[95] (e.g., the nontrivial $\mathbb{Z}$ class[107]) which suppress the short-range entanglement and thus promote the transition from condensated phase (topologically trivial phase) to the topologically ordered pure state (the product state representing an ensemble). In the case of strong Markovian noise, even the shortest-range (nearest neighbor) incoherent hopping in perturbative treatment would leads to diffusion[120]. Note that, different to the "Hermiticity" of $mm$-case where there is power law decay (also including the Gaussian decay of decoherence by the strong system-bath coupling[115]), the "Hermiticity" in late-stage of $sq$ case exhibits exponential decay where the locality is guaranteed by the large dissipative gap under nonunitary (disentangling) measurements[63, 104, 133], and this corresponds to the macroscopic response to the quantum jumps measured in a timescale (waiting time) larger than the fluctuation (scattering induced by imaginary potential[122]) timescale. Equivalently, the fast relaxation to the (pure) dark state in $sq$-case contributes to the global area law phase (e.g., the $Z_{2}$-type localized phase)[63, 15]. The condensed phase here in $mm$-case does not involve the fluctuation and dissipation stochastic contributions from the quantum jumps, and the power-law decay guarantee the (solvable) nonperturbative characteristic[104, 116, 117, 127] (until the sparse measurements and dissipations as well as perturbative contribution from jumps set in). Despite as a hallmark of valid bosonization and condensation and kinetic constraint in Hilbert space[127, 104, 51], the diffusive power-law scaling of long-range hopping and fluctuations is available in $sq$-case (attributed to the ergodicity and weak measurements) [56, 126, 39, 96, 128, 129, 130], e.g., the system with anisotropic power-law long-range dipolar interactions[132, 134, 142, 153] which contributes to the polarization and conservation of collective spin and critically slow thermalization (e.g., slow relaxation of magnetization or spin polarization)[142, 153, 145] in contrast to the fast Markovian dephasing (e.g., the global depolarizing noise channel[174, 175]).
Distinct from our above statement in many-body perspective that the $mm$-case is dominated by algebraic decay with power-law short-range correlations (due to the skin effect induced, e.g., by the sparse measurements) and the depinning of particle density caused by unitary dynamics (delocalized phase[15]), each single-particle Gaussian state in $mm$-case exhibits exponentially decaying short-range correlation (with area law entanglement) in the case of logarithmically growing (unbound) correlation length[63], e.g., in massive boson mode, but such area law phase induced by exponentially decaying correlations (together with the macroscopic wave function for photons or condensed bosons, i.e., paired fermions with short-range hopping) only takes a little fraction in the $mm$-case due to the isotropic nature and the encoded robust initial state information as long as the system is dominated by the condensation. This is also in agreement with [136], where the classical stochastic dynamics can be overwhelmed from the many-body microscopic details. Despite the skin effect, $mm$-case is dominated by the algebra (power-law) decaying, and indefinitely heating up toward the infinite temperature and maximally mixed steady state, as can be evidenced by the absence of a characteristic time scale in the long-time dynamics and the presence of significant phase fluctuation (suppressed by the sparse measurements and implies no thermalization[134, 135, 136, 137]; see Fig.2) and gapless structure of Hermitian measurement operator or Liouville superoperator[133]. Thus, the entanglement entropy loss (heat loss) induced by the sparse quantum jumps is overwhelmed by the entanglement entropy gain due to the non-Hermitian evolution in each sector as well as the non-Gaussian interactions[136, 134].
Thus in perspective of the whole system, the topological order pure state with long-range entanglement and lower measurement rate is favored in $sq$-case where particle-hole symmetry is broken (irreversible) by the long-range hopping, while disordered phase in constricted Hilberts space with short-range entanglement and higher measurement rate is favored in $mm$-case where particle-hole symmetry is preserved (reversible) by the short-range hopping[104, 122, 106, 107]. The topological nontrivial phase in $sq$-case, as stabilized by engineered dissipations, refers to the global symmetries, like the $\mathbb{Z}$ class, to guarantees the bulk-boundary correspondence or the valid description of winding number; While the topological trivial phase refers to the discrete symmetrie, like the particle-hole symmetry. Such disordered-to-ordered transition is also reminiscent of the strong-to-weak spontaneous symmetry breaking (in opposite to the effect of local decoherence from a short-depth local quantum channel[81, 78, 44, 113, 114, 117, 153]), or the reversible-to-irreversible transition, where a description of bosonization is valid in Hermitian limit ($mm$) as evidenced by the linear infinite temperature average (Gaussian correlation) as well as the unsuppressed heating up and noise-induced fluctuations[104, 112, 63, 116, 121], but breaks down in low-density regime with nonunitary dynamics and nonperturbative treatment under continuous post-selected measurements[104, 63]. The Anderson localization is favored under stronger (static) disorder in the late-time stage with saturated localization length. Thus the $sq$-case with Wiener process (Brownian motion)[97, 98, 90] involving both the deterministic (single trajectory level) and Ito stochastic under weak (dephasing) noise. While a unitary dynamics emerges when the uniformly measurement is performed to ensure constant dissipation rate $\gamma_{i}$, and $\sum_{i}L_{i}^{{\dagger}}L_{i}$ is bounded to be a conserved quantity which is inversely proportional to the waiting time between projective measurements[90].
For weighted trajectories in $mm$ case, the unitary evolution guarantees the open boundary condition, and the well preserved initial quantum information due to the interacting effect (in a non-Hermitian way) is consistent with the periodic drive which is widely seen in realizing the unitary model, as a example, the ensemble randomness of unitary gates in a unitary circuit induced the chaos (short-range entanglement). Thus the local unitary gates equivalents to the effect of nonlocal measurement, reminiscent of the forward scattering of thermal eigenstates. In classical configuration, the randomness in locations of projective measurement is necessay to observe volume-law to area-law phase transition through the zeroth Rényi entropy $S_{0}={\rm ln}\ {\rm Rank}(\rho)$, which is sensitive to the jump-induced weak perturbaation due to the rank-dependence[83] and has less contriution to the entropy[66]. where the ratio $S_{0}/{\rm ln}t$ exhibits scale invariance at critical measurement rate with the logarithmic long-range entanglement[83, 44, 84] This is consistent with the classical anology of the measurement-induced entanglement phase transition, like the local-projection-induced information freezing[86], or the driven-induced synchronization[85, 109].
Figure 1: Von Neumann entropy $-{\rm ln}{\rm Tr}D^{2}$ for $sq$ (left) and $mm$ (right) cases, which can be termed as microscopically defined entropy and thermodynamic entropy, respectively. Blue, orange to green dots represent the increasing system size at balanced partitions. The transition of linear (volume law) scaling to saturation toward a system-size-independent value (area law) can be seem in $sq$-case. The linear growth of entropy in early time is consistent with the scrambling (spreading of local information) with constant slope described by classical Lyapunov exponent[160], Thus for different system size, there are different scrambling time but same Lyapunov exponent. while the quantum mechanics with interacting effect only sets in latter as exhibited by the power-law growth toward the maximal entropy which corresponds to the dip in purity (see Fig.7). Note that the maximal entropy here does not corresponds to the maximal mixture as shown in Fig.7. The entropy in $mm$-case is dominated by the quantum mechanics. The revival behavior in $mm$-case results in a periodically fluctuating entropy, which should be distinguished from the macroscopic mechanical oscillations[180] and can be fitted by the conformal field theory (discontinuous) in the analysis of finite size scaling[39, 149, 166]. We also verify that the entropy here equals to the negative of the logarithmic negativity $-{\rm ln}({\rm Tr}D^{1/2})^{2}$[155]. The scheme (a) automatically arrives the balanced partition due to the densest measurement; While the scheme (b) admits arbitary partitions as revealed by the entanglement between the unitarily or nonunitarily evolved pure state $|\rho_{0}\rangle\langle\rho_{0}|$ and the rest. For (a), it is reasonable to suspect that, the peak of entropy between the (sub)volume phase in early time and area law phase in late time, which sharpens with increasing size, will vanish in the large system size limit, and the scrambling will also vanish.
Figure 2: Evolutions of trace and norm of the dissipative term for $sq$ (blue) and $mm$ (orange) cases.. The exponential decay and incoherence feature can be seem from the $sq$-case, reflecting that the dissipative terms is nondegenerated or nondefectively degenerated. While the algebraic decay is hard to recognized from the $mm$-case, due to the complicated coherence (averaged) pattern. The left panel is indeed a conditioned expectation of the stochastic object which does not depend on Wiener increment (Markovian noise) but on stochastic increments at earlier times[119].
Figure 3: Loschmidt echo for $sq$ (left) and $mm$ (right) cases. First, second, and last rows corresponds to the full, nondegenerated, and degenerated (NDD) eigenspaces, respectvely.
Figure 4: (a) and (b) show the $e^{i(\rho_{i}+\rho_{j})t}-e^{i\rho_{i}t}e^{i\rho_{j}t}$ ($\neq 0$ only for $\rho^{{\dagger}}=\rho$) representing the periodically oscilatting classical correlations (or noncommutativity) between the two arbitary states of the $mm$-case, for incomplete and complete basis, respectively. Such pairwise correlations is unique to the $mm$-case, due to the lacking of macroscopic oscillation[180]. Blue, orange, and green dots represent the real, imaginary, and absolute parts. Such correlations, which exist between arbitarily two components in the $mm$-case, cause the delocalization and the weakened boundary driving effect. This also make the construction of Floquet driving potential available, and such correlation (in constrast to the entanglement) can also be reflected by the separability across all bipartitions[180, 182], which is realized by the projective measurements on one subsystem and cause the statistical mixture of pure states (chaotic projected ensemble) in the another subsystem (disentangling)[181]. Coherences are allowed between these mixed subsystems. (c) and (d) show the trace distace $\frac{1}{2}||D_{i}-D_{j}||$ (|||||$\cdot$||| | ⋅ | | representing the trace norm), which witness the distinguishability between states $D_{i}$ and $D_{j}$ (initially pure states), for incomplete and complete basis, respectively, with blue and orange dots correspond to the $sq$-case (blue) and $mm$-case (orange), respectively. For complete basis, a uncontinuously reducution of trace distance for $sq$-case reveals that the existence of non-Markovian process in early stage. Despite not being shown here, we have verified that, the Jensen-Shannon divergence[179] of $sq$-case reduces to zero at long-time. Still, the non-Markovianity in $mm$-case, as revealed by the memory effect and environment-to-system information backflow, causes the nonmonotonic behavior of trace distance as well as the distinguishability between the two quantum states, and a seperable mode into the hot and cool ones. Similar nonmonotonic behavior in entanglement and decoherence function are studied in Refs.[178, 183]. The trace distance in (d) for $mm$-case exhibits persistent temporal fluctuations despite more and more homogeneous compares to initial as indicated by the dashed line. It is the correlations cause the revival of trace distance and environment-to-system backflow of information. Ref.[173] indicates a possible realization of such a scenario by the balanced interplay between interaction (with bath) and tunneling.
Figure 5: Von Neumann entropy for product states $(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}^{*}|)$ ($sq$) and $(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}|)$ ($mm$). The fixed bulk pattern here indicates the possibility to be free from thermaliation, which disappears in $sq$-case with the nonunitary evolution which varies the upper and lower boundaries. Similar to Fig.1(a), a gapless (volume law with density depinning) to gapped (area law with density pinning) transition can be seem for $sq$-case (i.e., nonequilibrium boundary to equilibrium boundary crossover[157] with exponentially slow disentangling consisting with Lorentzian distribution). For product state in $mm$-case, even the local thermalization (induced by the coupling to degrees of freedom of bath) cannot be seen. The fully engineered entropy evolution in long-time (by the dissipative environment) in $sq$-case reflects the difference between the few-body cluster entanglement and the many-body (fluctuating) one.
Figure 6: (Left) $|{\rm Tr}[\frac{D_{mm}(t)}{||D_{mm}(t)||}]|$ (blue), $|{\rm Tr}[\frac{e^{-i\rho_{mm}t}}{N||e^{-i\rho_{mm}t}||}]|$ (orange), $|{\rm Tr}[\frac{D_{sq}(t)}{||D_{sq}(t)||}]|$ (green), and $|{\rm Tr}[\frac{e^{-i\rho_{sq}t}}{N||e^{-i\rho_{sq}t}||}]|$ (carmine). The last two approach to $\sqrt{\frac{1}{N}}$ and $\frac{1}{N}$, respectively, in long-time. The variance of non-Hermitia observable[180] as well as the Jensen-Shannon divergence (JSD) also support the conclusion that the $mm$-case has much robust uncertainty compares to the $sq$-case. Note that here the uncertainty and randomness are not due to the selections by measurement apparatus, which mean the Green function here does not contains a self-energy term. (Right) The expectation $|{\rm Tr}[\frac{D_{sq}(t)}{||D_{sq}(t)||}]|=\langle\rho_{0}|e^{i\rho_{sq}t}|\rho_{0}\rangle$ (Orange; without consider the perturbation from measurement apparatus) and $\langle\rho^{\prime}_{0}|e^{i\rho_{sq}t}|\rho^{\prime}_{0}\rangle$ (Blue; consider the perturbation from measurement apparatus). The effect measurement apparatus here is small but finite due to the nonunitarility of $e^{i\rho_{sq}t}$. The traditional mass perturbation theory is invalid here. The postselection probability after measurement $|\langle\rho^{\prime}_{0}|\rho_{0}\rangle|^{2}$ (in the absence of interaction) is shown in color green. While in $mm$-case, the projective measurement suppress the long-range entanglement but does not cause the selections on measured states due to the effect of unique Markovian bath. Another reason is that we choose a time-independent basis $\langle\rho_{0}|\rho_{0}\rangle$. Consequently, the mapping $|\rho_{0}\rangle\rightarrow|\rho^{\prime}_{0}\rangle$ does not cause any new conservations. This indeed reflects the existence of response to the global quench in $mm$-case, just like the complete thermal state in late stage of $sq$-case, i.e., the common characteristic between the Markovian system in $sq$-case and the unique Markovian bath in $mm$-case.
Figure 7: Purities ${\rm Tr}|D|^{2}$ for $D=D_{sq}$ (a) and $D=D_{mm}$ (b). Despite the global symmetry in $sq$-case, the purification back to the pure state is only available in the $mm$-case. This is because, in $sq$-case, the non-Markovian bath plays an essential role in engineering the dynamics of the system. The inset in (a) shows the entropy (the same with Fig.1(a)) for a comparasion. There are two times of entanglement enhancement (purity decaying) in $sq$-case, which are due to the scrambling and the combination of monitoring and boundary-driven effect, respectively. The red arrows in (a) indicate the time when the long-range correlation as well as the area-law scaling become dominating (entanglement becomes nonextensive due to the absence of scrambling), implying the exponential decay and strong localization[159] such that the Gaussian (exponentiated bilinear) is achievable. In late-stage of $sq$-case (after the red arrows), the continuous monitorng from the dissipative environment cause the decay of entropy to the finite value.
V Method and Numerics
We define the densities of the two classes as $\rho_{mm}:=\sum_{i}|\psi_{i}\rangle\langle\psi_{i}|$, $\rho_{sq}:=\sum_{i}|\psi_{i}\rangle\langle\psi^{*}_{i}|$. For $\rho_{sq}$ such asymmetry construction (off-diagonal) of the density is also consistent with that in the content of Dicke states[169, 171]. The reduced form can be seem by the state of the combined system environment $|\rho_{SE}\rangle=\sum_{i}|\psi_{i}\rangle|\mathcal{E}_{i}\rangle$ with $\langle\mathcal{E}_{i}|\mathcal{E}_{j}\rangle=\langle\mathcal{E}_{i}|\mathcal{E}_{j}^{*}\rangle=\delta_{ij}$. The complex eigenstates are prepared aimed to obtaining the non-Hermitian operator ${\bf\Psi}$ whose elements follow the distribution of Gaussian ensemble, upto second and fourth moments, and reads, when consider two real eigenvalues, ${\bf\Psi}=\lambda_{\alpha}{\bf\Psi}_{\alpha}+\lambda_{\beta}{\bf\Psi}_{\beta}$ For a general route of construction, see Refs.[150, 8]. The densities $D_{mm/sq}(t)=e^{i\rho_{mm/sq}t}(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}|)$, whose square of trace is the Loschmidt echo, is important in revealing the difference between $sq$ and $mm$ cases. The initial state for both cases are pure state and unentangled. Available convex sum form for densities in $mm$-case (diagonal in the Schmidt basis[156]) agrees with the absence of topological order therein. Environmental decoherence effect (off-diagonal fluctuation) emerges when $\langle\mathcal{E}_{i}|\mathcal{E}_{j}\rangle\neq 0$ for $i\neq j$, and this will suppress the entanglement and cause the gapless structure for $mm$-case and early stage of $sq$-case. Such decoherence and off-diagonal fluctuation will be suppressed by larger system size in $sq$-case, and decay with time toward the gapped (area law). Also, the entropy for product state $(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}^{*}|)$ converges to $1/2^{N}\approx 0.0039$ (maximally mixed density) where $N=8$ here is the maximal number of independent states, which indicating the possible relation to the Page value[161]. Also, as shown in Fig.4, the correlation dynamics revealed by the expectation of covariance matrix (nonlocal density; fermion Green’s function)[164, 129, 170] ${\rm Tr}[D_{mm/sq}(t)]$, reflects that in $sq$-case a trajectory in a $2^{N}$-dimensional space can be encoded into the evolution of the $N\times N$ observable $e^{-i\rho_{mm/sq}t}$, where $\overline{\rho_{sq/mm}}=\overline{e^{-i\rho_{sq/mm}t}}_{t=0}=\lim_{t\rightarrow\infty}|{\rm Tr}[\frac{e^{-i\rho_{sq}t}}{N||e^{-i\rho_{sq}t}||}|=(\lim_{t\rightarrow\infty}|{\rm Tr}[\frac{D_{sq}(t)}{||D_{sq}(t)||}]|)^{2}=\frac{1}{N}$. A polynomial complexity in system size is only available in the late-stage of $sq$-case, where the simplification is available due to the absence of persistive fluctuates in Hilbert space[166] and the Lieb-Robinson bounds and nonunitary scrambling light cones. Also, the incoherent coupling to the environment is absent in the late-stage of $sq$-case which is in agree with Refs.[164, 129] . While such decoherence as well as stochastic dynamics is persistent in $mm$-case, due to the restricted Hilbert space and disorders which guarantee the accessibility of (maximally mixed) infinite-temperature featureless equilibrium state in long-time limit. Despite the classical stochastic processes, like the random diffusion, still observable in $mm$-case, the dynamics periodically back to the initial state, consistent with the microscopic theory. Such recurrence is recently studied in terms of the uncertainty relation[163, 165] and hybrid quantum circuit[166].
[…]
Consistent with the results in Ref.[184], here the non-Hermitian and unitary observables $O_{nH}=e^{-i\rho_{mm}t}$ ($t\neq 0$) is full-rank such that it can be decomposed into $O_{nH}=U_{mm}R_{mm}$ with positive definite $R_{mm}=\sqrt{O_{nH}^{{\dagger}}O_{nH}}$ and $U_{mm}=\frac{O_{nH}}{\sqrt{O^{{\dagger}}_{nH}O_{nH}}}$ ($U_{mm}^{{\dagger}}U_{mm}=I$). Then we can verify that $\langle\rho_{0}|O_{nH}|\rho_{0}\rangle=\langle\rho^{\prime}_{0}|R_{mm}|\rho_{0}\rangle=\langle\rho^{\prime}_{0}|O_{nH}|\rho^{\prime}_{0}\rangle$, with $|\rho^{\prime}_{0}\rangle=U_{mm}^{{\dagger}}|\rho_{0}\rangle$ and $\langle\rho^{\prime}_{0}|R_{mm}|\rho_{0}\rangle/\langle\rho^{\prime}_{0}|\rho_{0}\rangle=1$. The role of measurement apparatus can then be taken into account by preselecting the $|\rho_{0}\rangle$, or equivalently, postselecting the $|\rho^{\prime}_{0}\rangle$, where the perturbations from measurement can be revealed by the nonzero imaginary part ${\rm Im}\langle\rho^{\prime}_{0}|R_{mm}|\rho_{0}\rangle/\langle\rho^{\prime}_{0}|\rho_{0}\rangle\neq 0$. Thus the mapping $|\rho_{0}\rangle\rightarrow|\rho^{\prime}_{0}\rangle$ does not generates any new symmetries, instead, zero imaginary part of $\langle\rho^{\prime}_{0}|R_{mm}|\rho_{0}\rangle/\langle\rho^{\prime}_{0}|\rho_{0}\rangle$. guarantees the existence of response to the global quench in $mm$-case.
In the numerics as shown in Fig.2, we consider the case totally dominated by the dissipative process such that the unitary part (i.e., evolution and decoherence due to the system Hamiltonian) in Lindblad master equation is vanished. This cause the complete fail of perturbative treatment (e.g., the one introduced in Ref.[133] which is implemented by considering the steady state of nonperturbed part be a pure state).
An efficient way to enhance the noise from the environment’s measurement is replacing the $-{\rm ln}{\rm Tr}D^{2}$ by $-{\rm ln}\sum_{i}{\rm Tr}\rho_{i}^{2}$ where $\sum_{i}\rho_{i}=D$, which cause a long-time steady state closer to a product state with much lower entanglement compares to Fig.1 and similar in mechanism to Fig.LABEL:FK (by fill in the off-diagonal components such that it becomes fail to estimate the expectation value from local observable like the single-particle macroscopic sample). Obviously, this also corresponds to a purifying for $sq$-case.
For Loschmidt echo shown in Fig.3, we define the initial state $|\rho_{0}\rangle=\sum_{i}|\psi_{i}\rangle$, and estimate the overlap $\mathcal{L}(\rho,\rho^{\prime},\cdots):=\frac{|\langle\rho_{0}|e^{-i\rho t}e^{i\rho^{\prime}t}\cdots|\rho_{0}\rangle|^{2}}{||e^{i\rho t}e^{-i\rho^{\prime}t}\cdots||^{2}}$. Note that for $sq$-case, due to the nonunitarility, a artifical renormalizaton by the squared norm is necessary to avoid the diverging as we shown in Ref.[147]. In nondegenerated region, eigenstates in $sq$-case are mutually independent, while those in $mm$-case are mutually dependent. This can be easily verified by comparing $\mathcal{L}(\rho+\rho^{\prime},\cdots)$ and $\mathcal{L}(\rho,\rho^{\prime},\cdots)$, e.g., $\frac{|\langle\rho_{0}|e^{-i\rho_{mm}t}|\rho_{0}\rangle|^{2}}{||e^{-i\rho_{mm}t}||^{2}}$ and $\frac{|\langle\rho_{0}|e^{-i\sum_{i=1}^{j}|\psi_{i}\rangle\langle\psi_{i}|t}e^{-i\sum_{i=j+1}^{N}|\psi_{i}\rangle\langle\psi_{i}|t}|\rho_{0}\rangle|^{2}}{||e^{-i\sum_{i=1}^{j}|\psi_{i}\rangle\langle\psi_{i}|t}e^{-i\sum_{i=j+1}^{N}|\psi_{i}\rangle\langle\psi_{i}|t}||^{2}}$. While in NDD region, each eigenstate therein is independent of arbitary other one, producing the long range by system-size-independent correlations. The most significant feature of this region is the two parallel by seperated boundaries in echo spectrum. This is a special region where the gap is not supported by coherences and will not prohibit the global thermalization (when there are system-size-dependent correlations).
VI Conclusion
The steady state of $mm$-case (consistent with the open boundary condition) is only due to the skin effect, where the many-body state (without exponential localization) holds global fluctuation due to the Hermiticity and the long-time behavior can be well predicted by diagonalization. This is in contrast to the localized state in $sq$-case (consistent with the periodic boundary condition) where the single-particle eigenstates exponentially decay with the localization length independent of system size[56, 94, 10, 124, 125, 130], and thus the long-time behavior is harder to predict by approximately diagonalization due to the non-Hermiticity. Zeno effect and confined light cone[131, 96] are unique to the $sq$ case, while Hilbert space fragmentation[154, 121] and linearized bosonization are unique to the $mm$ case, and the topological islands[125] is unique to the intermedia regime with finite strength of disorder but not yet reaches the ensemble-averaged result. A similar study with the correspondence to quantum systems is available by by, e.g., transverse-field Ising model[162], XXX spin model, or BoseHubbard model[161], where the ergodicity (late stage) and scrambling (early stage) of $sq$-case are shared by the qubit ladder and nonintegrable transverse-field Ising model, while the nonergodicity of $mm$-case is shared by the qubit chain and integrable XXX spin model. Also, a protocol similar to the continuous measurement which give rise to the diffusion is the collective jumps[168] where the collectiveness can be diagnosed by the enhanced entanglement. Intermediate cases is achievable through complex disorder patterns by turning the transverse and longitudinal magnetic fields where the scrambling occurs with temporal fluctuations suppressed by the disorders (ensemble average) and initial state becomes mixed up in a large subspace[162]. While the absence of scrambling in $mm$-case is due to the conservation of initial states and each subspace (block-diagonal unitary operator) owns a unique exponentially decaying behavior, anology to the internal dynamics of black hole and local integrals of motion of many-body localization which suppress the scrambling and rise the fluctuations. Note that the constriction on Hilbert space of the $mm$-case should be distinguished from the restriction of Hilbert space in $sq$-case where for the latter it is implemented by the construction of, e.g., Dicke basis (classical mixture cause thermalization as guaranteed by the permutational symmetry) and reducing from the $O(2^{N})$ Hilbert space dimension to $O(N)$ subspaces.
Data Availability Statement: The data that support the findings of this study are available within the article.
VII Appendix: Stochastic evolution with Markovian and non-Markovian dynamics
The numerical operator $\sum_{i}L^{{\dagger}}_{i}L_{i}$ is commutes with the nondissipative (Hermitian system) Hamiltonian. Intriguing phenomenas, like superradiant burst[29] during the photon emission, arise by choosing a Hermitian dissipation or decoherence matrix ${\bm{\gamma}}$ with nonzero off-diagonal component. Note that this dissipation strength is related to local imaginary potential[108, 122, 104] (inversely proportional to the pulsed measurement period), and is also referred to as measurement rate (or measurement strength), or the non-Markovian/Markovian noise amplitude, and proportional to the system-bath coupling strength[139].
The system of $sq$-case in Born-Markovian approximation weakly couple to the non-Markovian thermalizing environments where the non-Markovian noise therein cause the dephasing and broadened distribution with long-range hopping. Also, the memory effect of non-Markovian environments contributes to the large dissipation and the preserving of entangled states within the system as well as the long-time destination to larger steady-state entanglement (compares to the Markovian environments where the Markovian noise mainly cause the fast fluctuations)[120, 116].
For non-Markovian dynamics non-local in time, the dynamics can nolonger be described by Lindblad master equation and the corresponding stochastic Schrodinger equations can describes only the no-jump process (controlled dissipation), instead of the jump (unconditioned dissipations) upon a measurement record[34, 167]. There is finite detector bandwidth and noise-noise correlation due to the information backflows from environment to system (in contrast to the Markovian system with nonunitary Lindblad dynamics[116]), despite the description by Markovian master equation still hold on average[34] A vanishingly small dissipative incoherent case $\gamma\rightarrow 0$ is applied to non-Markovian dynamics to guarantees the fast relaxation and dependence on noise correlation time, in the absence of quantum jump [90, 80, 96, 102]. Thus, despite the coupling to the non-Markovian environments, the continuous measurement by homodyne or heterodyne detection can be realized in $sq$-case in Markovian limit (i.e., infinite bandwidth, infinite count rate, and perfect detector), where the large dissipation rate guarantees the thermal state and the loss of initial information due to the environment. In $sq$-case, the volume-law (saturated) entanglement in short-time (long-time) scale is dominated by the Markovian (non-Markovian) dynamics.
Differently, in $mm$-case, evolution of the system strongly coupled to a thermal reservoirs (Markovian environment) is well described by the Markovian Lindblad master equation, and the non-Markovian memory effect induced complex coherence pattern (the infinite bound states due to the dissipationless non-Markovian dynamics), whose backaction effect induce the oscillation in entanglement, and the steady state is possible only in long-time limit. Note that despite the strong system-bath coupling, the Markovian thermal environment equilibrium at a certain temperature initially coherent with the system (without the quantum correlation/entanglement) in a product form $\rho_{S}\otimes\rho_{E}$[102, 103]. In this case, the strong noise in Markovian (noisy) environments results in a long timescale of slow relaxation, and the long non-Markovian noise-noise correlation time (weak non-Markovian) within bath in fact has a indistinguishable effect with the Markovian white noise, where both give rise to the classical diffusion in the system. In interacting case, the dissipative interaction is represented by a decoherence matrix which push the system toward thermalization, In noninteracting limit, ${\bm{\gamma}}_{0}=\gamma{\bf I}$ while in the all-to-all interacting limit (Dicke limit in canonical setting[33, 32, 168]), ${\bm{\gamma}}_{D}=|\sqrt{\gamma}\rangle\langle\sqrt{\gamma}|$, with $|\sqrt{\gamma}\rangle$ the vector consist only the element $\sqrt{\gamma}$. The vectors made of eigenvalues arranged in descending order for ${\bm{\gamma}}_{D}$ supermajorize that of ${\bm{\gamma}}_{0}$.
In Markov limit where the equation of motion depends only on the nearest past time, the bath energy density becomes static and depends only on the dissipation rate, while the general non-Markovian processes involve the average over the whole past (the noise),. In strong non-Markovianity limit, the bandwidth of bath spectral density distribution is vanishing small and results in band-gap structure in bath, where the dissipation (as well as the system-bath coupling) favors the coherent dynamics[73, 47]. For systems of non-Hermitian dynamics, like the $mm$-case, the superpositional coherences coexist with the nonequilibrium dynamics induced by system-bath coupling, and the bound states preclude the exponential decay as in single particle picture. Such nonequilibrium dynamics in a open system the delocalization coexist with the coherence-induced many-body dynamical localization contribute to the large fidelity and the stored quantum information that is robust against disorder (thermal noise).
In the absent of quantum jump (by postselection), the dynamics can be described by the Linbladian equation $\frac{d}{dt}\rho=-i[H,\rho]-\sum_{i}\frac{\gamma_{i}}{2}\{L^{{\dagger}}_{i}L_{i},\rho\}=-i(H_{{\rm eff}}\rho-\rho H^{{\dagger}}_{{\rm eff}})$ with $H_{{\rm eff}}=H-i\sum_{i}\frac{\gamma_{i}}{2}L^{{\dagger}}_{i}L_{i}$, where the imaginary part of no-jump measurement operator $L_{0}=(\gamma_{0}dt)^{-1/2}(1-iH_{{\rm eff}}dt)$ only given by the Hamiltonian $H$ which governs evolution of $\rho$. The measurement-induced back action $i\sum_{i}\frac{\gamma_{i}}{2}L^{{\dagger}}_{i}L_{i}$ is due to the memory effect non-Markovian environment[139, 151] which may contains some unobservable degrees of freedom coupled to the system and cause the thermlization[151]. The Lindblad master equation consistent with the continuous limit of quantum channel[75, 60], or the Markovian equation of motion of density matrix[123]. The individual trajectories $|\psi(t)\rangle=\frac{e^{-iH_{{\rm eff}}t}|\psi_{0}\rangle}{||e^{-iH_{{\rm eff}}t}|\psi_{0}\rangle||}$ have the stochastic time evolution[34, 91] (expanded up to first order in infinitesimal measurement time $dt$ during the monitoring) $d|\psi(t)\rangle=(-iH-\sum_{i}\frac{\gamma_{i}}{2}L^{{\dagger}}_{i}L_{i}+\langle\psi(t)|\sum_{i}\frac{\gamma_{i}}{2}L^{{\dagger}}_{i}L_{i}|\psi(t)\rangle)|\psi(t)\rangle dt$ in the absence of detection of (local Poisson) quantum jump (photo emission). The existence of time-dependent term $\langle\psi(t)|\sum_{i}\frac{\gamma_{i}}{2}L^{{\dagger}}_{i}L_{i}|\psi(t)\rangle$ within above equation is due to the time-dependent norm $||e^{-iH_{{\rm eff}}t}|\psi_{0}\rangle||$ where $L^{{\dagger}}_{i}L_{i}$ and $H$ are Hermitian but $H_{{\rm eff}}$ is non-Hermitian. In this case the measurement on a individual trajectory returns no result and the pure state $|\psi(t)\rangle$ changes infinitesimally under the non-unitary evolution. In above Lindblad equation, the commutator is the unitary part which is related to the coherence, while the summation over quantum channel index is the dissipative part. When the quantum channel index also refers to the local site index, the system-bath coupling on these local sites will causes the exponential decaying in Lieb-Robinson velocity of local observable[72], and leads to exact ground-state in infinite time limit[82]. Such slow down of quantum information propagation in the delocalization (chaotic) phase is consistent with the effect of disorder[52, 86] and the exponential localization is in agree with the thermalization in single particle picture where the system-bath coupling in moderate strength results in a unique steady state without the memory of initial state information. While for strong system-bath coupling (dephasing), the rised purely classical dynamics will suppress the thermalization by diminishing the mutual information (quantum correlations)[72, 28]. Both the quantum information which propagating in fixed velocity against the effect of disorder, and the coherence in open many-body system coupled to Markovian bath , exhibit nonlocal nature, which dominates in the non-Hermitian system with superpositional coherence pattern as a mixed state (as long as there are more than one trajectory) whose trajectory-averaged time evolution of the density matrix under Lindblad master equation is conditioned to a certain number of quantum jumps between trajectories (a special case with null jumps is the non-Hermitian case where there is a single trajectory)[75, 60, 76, 39]. The nonlocal measurement applied on such nonequilibrium many-body system reveals the incompatibility with the single-particle picture (i.e., the no-click limit with rigorous bulk-edge correspondence) or the mean-field non-Hermitian dynamics where dissipative and decoherence dominate. The probe record of such quantum jumps in unfixed number is not known. This is consistent with the $mm$-case, where there are infinite bound states corresponding to the mode propagating in growing speed and not be bounded by Lieb-Robinson velocity (the exponentially decaying). Such information propagating could be carried by supersonic mode[75], or the fluctuation-suppressed quasiparticle mode[10].
For discontinuous measurement protocol, distinct to the Homodyne or Heterodyne continuous measurement, there are discrete countable conserved quantities, like photon, and the corresponding unconditional dissipative dynamics is described by the Lindblad master equation $\frac{d}{dt}\rho=-i(H_{{\rm eff}}\rho-\rho H^{{\dagger}}_{{\rm eff}})+\sum_{i}\gamma_{i}L_{i}\rho L_{i}^{{\dagger}}$. For this protocol, despite there are allowed transitions between nonunitary and unitary evolutions by the absence and presence of quantum jump which return null and unit detection results (Poisson random stochastic increments), respectively, during an infinitesimal time interval, the system only exhibit jump evolution and never enters the regime of diffusive evolution, as described by the point-process stochastic Schrodinger equation (nonlinear in $|\psi(t)\rangle$),
$d|\psi(t)\rangle$ $=|\psi(t+dt)\rangle-|\psi(t)\rangle$ (4)
$=-i\left(H_{{\rm eff}}+\frac{i}{2}\sum_{i}\gamma_{i}\langle\psi(t)|L^{{\dagger}}_{i}L_{i}|\psi(t)\rangle\right)|\psi(t)\rangle dt+\sum_{i}\left(\frac{L_{i}}{\sqrt{\langle\psi(t)|L^{{\dagger}}_{i}L_{i}|\psi(t)\rangle}}-1\right)|\psi(t)\rangle dN_{i}(t),$
where the expectation (or ensemble average) on the Poisson random stochastic increment $dN_{i}(t)$ is the Born probability for unit detection result, and is unknown until the measurements are done, which reads ${\mathbb{E}[dN_{i}(t)]}=\langle\psi(t)|L^{{\dagger}}_{i}L_{i}|\psi(t)\rangle\gamma_{i}dt=||L_{i}|\psi(t)\rangle||^{2}\gamma_{i}dt$, $dN_{i}(t)dN_{j}(t)=\delta_{ij}dN_{i}(t)$. This makes it unable to eliminate or multiplicate the noise[91, 45], and meanwhile, retains the Poissonian randomness, in constract to the Wiener process with Ito stochasticity. Up to large number limit where the homodyne scheme and quantum state diffusion are achieved, the Poisson $dN_{i}$ is replaced by the multiplicative Gaussian Ito noise. The local (short-range) correlations in single-particle picture (closed system) corresponds to the quantity linear on the stochastic quantum states like the expectation in basis of stochastic states, and the absence of long-range order is consistent with the result of cluster decomposition[79, 80].
While the nonlocal correlations corresponds to the nonlinear one, like the von Neumann entropy and von Neumann correlation. Only the latter is free from Lieb-Robinson bound and deeply affected by the local quantum channel, e.g., the ensemble average over a large number of quantum jumps such that the local quantum channel has a large enough width to take effect while a narrow quantum channel (mixed state) is in the sense that it can be purified to a symmetric local unitary operator[78], and local correlations are avoided by the statistical mixture which generally corresponds to weak symmetry[39, 77, 78]. Also, this is also in agree with the suppression on area-law entanglement by the quantum jumps which plays the role of perturbation. Further, in the case neither dominated by the classical dynamics or reaching the thermalization, the quantum information over large distance is allowed in inhomogeneous steady state with extensive long-range entanglement or the postquenched homogeneous integrable state[74].
VIII Appendix.B: Effect from Lindblad dynamics to $sq$-case
For non-Hermitian and nonunitary observables $O_{nH}=e^{-i\rho_{sq}t}$ ($t\neq 0$), $|\langle\rho_{0}|O_{nH}|\rho_{0}\rangle|=|\langle\rho^{\prime}_{0}|R_{sq}|\rho_{0}\rangle|\approx|\langle\rho^{\prime}_{0}|O_{nH}|\rho^{\prime}_{0}\rangle|$, and $\langle\rho^{\prime}_{0}|R_{sq}|\rho_{0}\rangle/\langle\rho^{\prime}_{0}|\rho_{0}\rangle\neq 1$, which means the effect of measurement apparatus keep finite but small during the evolution (also implies the perturbative treatment is invalid here). Then the effect of imaginary part related to the measurement-induced interaction can be estimated from the difference $\langle\rho_{0}|O_{nH}|\rho_{0}\rangle-\langle\rho_{0}|UO_{nH}U^{{\dagger}}|\rho_{0}\rangle$ where $U$ is nonunitary such that $[O_{nH},{\rm ln}U]\neq 0$ and ${\rm Tr}[UO_{nH}U^{{\dagger}}-O_{nH}]=0$.
Figure 8: Von Neumann entropy $S=-{\rm ln}{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]$ for different system sizes.
We show the von Neumann entropy $S=-{\rm ln}{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]$ in Fig.8. Compares to the von Neumann entropy shwon in Fig.1(a), we see that the the peak at intermediate time vanishes and the entropy become monotonic and smooth over the whole evolution.
Next we consider the effect from weak interaction and the one body loss $\gamma$, which is also the decoherence function, the dynamics of time-dependent density $R_{sq}$ can be described by the Lindblad function especially in long-time limit,
$\frac{dR_{sq}}{dt}=-i(H_{int}R_{sq}-R^{{\dagger}}_{sq}H^{{\dagger}}_{int})+\frac{\gamma}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)+\frac{\gamma}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right),$ (5)
where ${\bf I}$ is the $N\times N$ identity matrix, and the interacting Hamiltonian reads $H_{int}=g|\rho_{0}\rangle\langle\rho^{\prime}_{0}|+h.c.$. Then we numerically found that, the dynamics of expectation of $\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}+(|\rho_{0}\rangle\leftrightarrow|\rho^{\prime}_{0}\rangle)$ can be well described by
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\dot{R}_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}+(|\rho_{0}\rangle\leftrightarrow|\rho^{\prime}_{0}\rangle)$ (6)
$=$ $-\frac{ig{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}-\frac{ig{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}$
$-\frac{ig{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}-\frac{ig{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}$
$+\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{\gamma}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}+\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{\gamma}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}$
$+\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{\gamma}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}+\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{\gamma}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}.$
We have the following relations at arbitary $t$,
$\frac{-i{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=-i{\rm Tr}[(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]=2\ {\rm Im}\ {\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}],$ (7)
$\frac{-i{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=-i{\rm Tr}[(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]=2\ {\rm Im}\ {\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}],$
$\frac{-i{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=-i{\rm Tr}[(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]=2\ {\rm Im}\ {\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}],$
$\frac{-i{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=-i{\rm Tr}[(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]=2\ {\rm Im}\ {\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}].$
Note that $\ {\rm Im}\ {\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]=-\ {\rm Im}\ {\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]$.
At long-time where $U_{sq}$ is trace preserving and $1/{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|=1/{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|=1/N$, there is symmetry for the exchange between states before and after measurement,
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]({\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]-1)}=\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]({\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]-1)}=-\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=-\frac{1}{N},$ (8)
$\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{2{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{1}{2N},$
Note the following relations,
$\frac{{\rm Tr}[R_{sq}]}{N}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]}{N^{2}}$ (9)
$=-\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]({\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]-1)}=-\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]({\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]-1)}\sim e^{-t},$
$||\ |\rho_{0}\rangle\langle\rho_{0}|\ ||_{2}=||\ |\rho^{\prime}_{0}\rangle\langle\rho_{0}|\ ||_{2}=||\ |\rho_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}=||\ |\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}=\frac{1}{N}||\ |\rho_{0}\rangle\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|\ ||_{2}$
$={\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]={\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]=N\ (\forall\ t),$
as shown in Fig.9(a) where $\langle R_{sq}\rangle_{1}$, $\langle R_{sq}\rangle_{2}$, $\langle\mathcal{D}\rangle_{1}$ represent the second, third, and fourth term of the first expression of Eq.9. While $\langle R_{sq}\rangle_{3}:=\frac{1}{N}|{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]|$ and $\langle R_{sq}\rangle_{4}:=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]}$ are shown inFig.9(a)-(b). At long-time where $U_{sq}$ still be trace preserving,
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]}{||[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|]||_{2}}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{||[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]||_{2}}=\frac{1}{N},$ (10)
where the 2-norm and trace norm satisfy $||[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]||_{2}=2||[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]||_{tr}=N$.
Figure 9: (a) Blue, orange and green dots represent the expectations $\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}$, $\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}$ and $\bigg{|}\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]}{||[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|]||_{2}}\bigg{|}$, respectively. (b)Blue, orange and green dots represent the real part, imaginary part, and absolute value of $\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|]}$.
While at long-time limit $t\rightarrow\infty$ where $U_{sq}$ is nomore trace preserving and $1/{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|=1/N,1/{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|=0$, we have
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]({\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]-1)}=-\frac{1}{N},$ (11)
$\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{1}{N},$
$\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]({\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]-1)}=0,$
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\infty,$
To obtain a long-time-stable result, we restrict the dissipative part of the dynamics of $R_{sq}$ (Eq.(5)) to only containing the state before nonunitary evolution (measurement), i.e., $|\rho_{0}\rangle$. then we have
$\lim_{t\rightarrow\infty}\left[\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\dot{R}_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}+(|\rho_{0}\rangle\leftrightarrow|\rho^{\prime}_{0}\rangle)\right]=\gamma\frac{2-N}{N},$ (12)
which is nearly $-\gamma$ in the large-$N$ limit. The long-time limit of $\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\dot{R}_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}$ turns to be the same result. This result is consistent with Refs., such that the dissipative part of the dynamics will not generates time-dependence to the postselection probability at long-time limit, as long as the decoherence function is time-independent.
Next we consider the effect from Lindblad dynamics to the time evolution of von Neumann entropy $S=-{\rm ln}{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]$,
$\dot{S}$ $=\frac{-1}{{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]}{\rm Tr}[\frac{dR_{sq}}{dt}|\rho_{0}\rangle\langle\rho_{0}|R_{sq}|\rho_{0}\rangle\langle\rho_{0}|+R_{sq}|\rho_{0}\rangle\langle\rho_{0}|\frac{dR_{sq}}{dt}|\rho_{0}\rangle\langle\rho_{0}|]$ (13)
$=\frac{-2}{{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]}{\rm Tr}[\frac{dR_{sq}}{dt}|\rho_{0}\rangle\langle\rho_{0}|R_{sq}|\rho_{0}\rangle\langle\rho_{0}|].$
We found that the unitary part in the Lindblad function shows to be irrelavant to introduce large randomness determined by the interaction strength $g$, thus here we consider only the role of dissipative part, i.e., $g\ll\gamma$. Then we define
$f_{1}:={\rm Tr}[\frac{1}{2}(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\})|\rho_{0}\rangle\langle\rho_{0}|R_{sq}|\rho_{0}\rangle\langle\rho_{0}|],$ (14)
$f_{2}:={\rm Tr}[\frac{1}{2}(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\})|\rho_{0}\rangle\langle\rho_{0}|R_{sq}|\rho_{0}\rangle\langle\rho_{0}|],$
$f_{3}:={\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}].$
As shown in Fig.10, $f_{1}$ exhibits the same exponential decay with $f_{3}$, while $f_{2}$ decay with fluctuations, and $f_{2}$ is bound by $f_{1}$ from below. Further, we found that, by renormalizing the functions such that their initial value equals 1, we have $\tilde{f_{1}}=\tilde{f_{3}}$. Thus the $f_{1}$ does not contributions to the time-dependence of the entropy, while $f_{2}$ soly leads to fluctuations with decaying amplitude. Roughly speaking, $\langle\rho^{\prime}_{0}|\rho_{0}\rangle$ contributes to the fluctuation last for a long time (as long as $U_{sq}$ be trace preserving), while $\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle$ and $\langle\rho_{0}|\rho_{0}\rangle$ contribute to the exponential decay with time. To understand this, we consider the correlation function between the two states (before and after measurement)
$\mathcal{C}_{1}(t):={\rm Re}\left(\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}-\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}\right)={\rm Re}\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}-1,$ (15)
where $\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}=\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}=1$ and thus $\mathcal{C}_{1}(0)=N-1$. We found that, in the late-time stage,
$\mathcal{C}_{1}(t)\approx-2\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=-2\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}\sim\alpha{\rm sin}(\beta t),$ (16)
which can be fitted by a sine function as shown in Fig.9(c). Here $\alpha$ and $\beta$ are the fitting parameters. We also consider the following correlation functions in varianted form
$\mathcal{C}_{2}(t):={\rm Re}\left(\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{||\ |\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}}-\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{||\ |\rho_{0}\rangle\langle\rho_{0}|\ ||_{2}}\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{||\ |\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}}\right)\sim\frac{-1}{{\rm sin}(\alpha t+\beta)^{2}},$ (17)
$\mathcal{C}_{3}(t):={\rm Re}\left(\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}-\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{||\ |\rho_{0}\rangle\langle\rho_{0}|\ ||_{2}}\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{||\ |\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}}\right)\sim\frac{-t}{{\rm sin}(\alpha t+\beta)^{2}},$
which can be fitted by the inverse squared sine function as shown in Fig.9(d). Using the identity
$|{\rm Tr}[\cdot]|=\frac{||(\cdot)^{2}||_{2}}{||\cdot||_{2}},$ (18)
we can obtain following relations,
$||\ |\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}=||\ |\rho_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}\ |{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]|=N|{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]|=N|{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|]|,$ (19)
$\frac{{\rm Tr}[R_{sq}]}{||\ |\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}}\approx\frac{1}{N}\bigg{|}\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]}\bigg{|}=\frac{1}{N}\bigg{|}\left(\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|]}\right)^{*}\bigg{|}\approx\frac{1}{N^{2}}\frac{(||\ |\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2})^{2}}{||\ (|\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|)^{2}\ ||_{2}}$
we have
$\mathcal{C}_{2}(t)$ $=(\mathcal{C}_{1}(t)+1)\frac{{\rm Tr}[R_{sq}]}{||\ |\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}}-\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}$ (20)
$\approx\frac{1}{N}(\mathcal{C}_{1}(t)+1)\bigg{|}\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]}\bigg{|}-(\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]})^{2}\sim\frac{\partial^{2}{\rm ln}\mathcal{C}_{1}(t)}{\partial\mathcal{C}_{1}(t)^{2}},$
$\mathcal{C}_{3}(t)$ $\approx(\mathcal{C}_{2}(t)+(\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]})^{2})\frac{(||\ |\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2})^{2}}{||\ (|\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|)^{2}\ ||_{2}}\sim\mathcal{C}_{2}(t)t.$
Figure 10: The blue, orange and green dots represent the renormalized functions $\tilde{f}_{1}$, $\tilde{f}_{3}$, and $\widetilde{f_{1}+f_{2}}$, respectively. The inset show the unnormalized functions $f_{1}$ (blue) and $f_{2}$ (orange).
Next we consider the off-diagonal term of the population, in which case the dissipative part reads
$\frac{\gamma}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho_{0}\rangle}{\bf I}-\frac{\{|\rho^{\prime}_{0}\rangle\langle\rho_{0}|,R_{sq}\}}{\langle\rho_{0}|\rho^{\prime}_{0}\rangle}\right)+(|\rho_{0}\rangle\leftrightarrow|\rho^{\prime}_{0}\rangle).$ (21)
Then by estimating its expectation on the states before and after the measurement, we obtain
$\frac{{\rm Tr}[\frac{1}{2}(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho_{0}\rangle}{\bf I}-\frac{\{|\rho_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}}{\langle\rho^{\prime}_{0}|\rho_{0}\rangle})|\rho_{0}\rangle\langle\rho_{0}|]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{{\rm Tr}[\frac{1}{2}(2\frac{\langle\rho_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\frac{\{|\rho^{\prime}_{0}\rangle\langle\rho_{0}|,R_{sq}\}}{\langle\rho_{0}|\rho^{\prime}_{0}\rangle})|\rho_{0}\rangle\langle\rho_{0}|]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]},$ (22)
$\frac{{\rm Tr}[\frac{1}{2}(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho_{0}\rangle}{\bf I}-\frac{\{|\rho_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}}{\langle\rho^{\prime}_{0}|\rho_{0}\rangle})|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=\frac{{\rm Tr}[\frac{1}{2}(2\frac{\langle\rho_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\frac{\{|\rho^{\prime}_{0}\rangle\langle\rho_{0}|,R_{sq}\}}{\langle\rho_{0}|\rho^{\prime}_{0}\rangle})|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}.$
Further, the expectations on the states before and after the measurement converge to a stable difference as long as $U_{sq}$ is trace preserving,
$f_{4}:=$ $\frac{{\rm Tr}[(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho_{0}\rangle}{\bf I}-\frac{\{|\rho_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}}{\langle\rho^{\prime}_{0}|\rho_{0}\rangle})|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}-\frac{{\rm Tr}[(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho_{0}\rangle}{\bf I}-\frac{\{|\rho_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}}{\langle\rho^{\prime}_{0}|\rho_{0}\rangle})|\rho_{0}\rangle\langle\rho_{0}|]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}\rightarrow\frac{1}{2^{N}},$ (23)
which vanish in the large-$N$ limit.
Figure 11: Eq.(23).